# SALT: Parameter-Efficient Fine-Tuning via Singular Value Adaptation with Low-Rank Transformation

### v1

@@ Abstract
The complex nature of medical image segmentation calls for models that are specifically designed to capture detailed, domain-specific features. Large foundation models offer considerable flexibility, yet the cost of fine-tuning these models remains a significant barrier. Parameter-Efficient Fine-Tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA), efficiently update model weights with low-rank matrices but may suffer from underfitting when the chosen rank is insufficient to capture domain-specific nuances. Conversely, full-rank Singular Value Decomposition (SVD) based methods provide comprehensive updates by modifying all singular values, yet they often lack flexibility and exhibit variable performance across datasets. We propose SALT (Singular Value Adaptation with Low-Rank Transformation), a method that selectively adapts the most influential singular values using trainable scale and shift parameters while complementing this with a low-rank update for the remaining subspace. This hybrid approach harnesses the advantages of both LoRA and SVD, enabling effective adaptation without relying on increasing model size or depth. Evaluated on 5 challenging medical datasets, ranging from as few as 20 samples to 1000, SALT outperforms state-of-the-art PEFT (LoRA and SVD) by 2% to 5% in Dice with only 3.9% trainable parameters, demonstrating robust adaptation even in low-resource settings. The code for SALT is available at: https://github.com/BioMedIA-MBZUAI/SALT.
Keywords:
Singular Value Decomposition, Parameter-Efficient Fine-Tuning, LoRA, Foundation Models, Domain Adaptation.
@@ 1 Introduction
Medical image segmentation plays a pivotal role in modern healthcare by enabling precise delineation of anatomical structures, pathological regions, and functional tissues across diverse imaging modalities such as CT, MRI, X-ray, ultrasound, and endoscopy [11, 20]. While Deep Learning (DL) has revolutionized this field, with architectures like U-Net [20] and nnU-Net [8] achieving state-of-the-art (SOTA) performance through sophisticated designs and extensive data augmentation, these models often rely on large parameter counts, limiting their adaptability to new datasets or modalities. For instance, nnU-Net’s success hinges on dynamic architecture adjustments and aggressive augmentation pipelines, while DeepLabv3+ [3] employs computationally intensive components like spatial pyramid pooling. Recent advances in foundation models have introduced specialized tools for diverse computer vision tasks. CLIP [18] and ALIGN [9], trained on large-scale image-text datasets, excel at cross-modal understanding for zero-shot image classification and open-set retrieval. In contrast, the Segment Anything Model (SAM) [10, 19], trained on billions of natural images, exhibits robust zero-shot generalization capabilities and initially raised hopes for universal segmentation. However, SAM’s performance on medical data has proven suboptimal, with studies [26, 6] demonstrating its inferiority to specialized models like U-Net++ [27], particularly in Dice scores across 12 medical datasets. Even with prompt engineering, SAM’s zero-shot accuracy remains below fully supervised methods [4], underscoring the domain gap between natural and medical images.
@@
Recent efforts to adapt SAM to medical imaging, such as MedSAM [13], SAM Adapter [23], and S-SAM [16], have made strides by incorporating medical training data, low-rank adapters, or SVD-based tuning. Yet challenges persist: PEFT methods must balance minimal parameter overhead, computational feasibility, and preservation of pre-trained knowledge while addressing unique medical imaging challenges like speckle noise, low contrast, and modality-specific artifacts. Promising approaches that arose to adapt foundational models, especially language models, like LoRA [7] and AdaLoRA [25] approximate weight updates via low-rank matrices but struggle to capture both dominant and nuanced features, while SVD-based methods like PiSSA [15] focus on principal singular vectors to capture the most significant data patterns, but this rigid prioritization can limit their adaptability to dynamic or fine-grained data characteristics. To address these limitations, we propose Singular Value Adaptation with Low-Rank Transformation (SALT), a novel PEFT framework that synergizes the strengths of SVD and low-rank adaptation. SALT selectively scales critical singular values to preserve foundational knowledge while applying trainable low-rank transformations to residual components, enabling efficient adaptation to medical domains with minimal parameter overhead. Evaluated on five challenging medical datasets, with sample sizes ranging from 20 to 1000 and demanding a high level of detailed segmentation, SALT consistently outperforms LoRA and SVD, achieving a 2 to 5% improvement in Dice score while utilizing only 3.9% trainable parameters. Our contributions are threefold:
  • We introduce SALT, a hybrid PEFT method utilizing SVD-based dominant singular value adaptation with low-rank residual updates, achieving parameter-efficient domain adaptation.
@@
  • We demonstrate SALT’s utility through extensive experiments on diverse medical datasets from various modalities, showing consistent improvements over state-of-the-art PEFT methods.
@@ 2 Methodology
Problem Formulation. Let S denote SAM model with pre-trained parameters $\Theta$, which consists of an image encoder ($\mathcal{E}_{\mathrm{i}}$), a prompt encoder ($\mathcal{E}_{\mathrm{p}}$), a mask decoder ($\mathcal{D}_{\mathrm{m}}$). $\mathcal{E}_{\mathrm{i}}$ contains $L$ blocks, each comprising a multi-head attention module and a Multi-Layer Perceptron (MLP). For a given input image $\mathbf{i}$ and prompt embedding $\mathbf{p}$, the encoder $\mathcal{E}_{\mathrm{i}}$ produces intermediate feature maps $\{\mathbf{f}^{(l)}\}_{l=1}^{L}$. Concretely, in the $l$-th block, the main operations can be written as:
$q,k,v=W_{qkv}^{l}\,i+b_{qkv}^{l},\quad o_{n}=W_{M}^{l}\,i+b_{M}^{l},$ (1)
where $W_{qkv}^{l},W_{M}^{l}\in\mathbb{R}^{D\times K}$ and $b_{qkv}^{l},b_{M}^{l}\in\mathbb{R}^{D}$ are the attention/MLP weights and biases for block $l\in\{1,...,L\}$. While effective for natural images, the pre-trained weights $W\in\mathbb{R}^{D\times K}$ may not capture all domain-specific nuances.
2.0.1 SALT.
Through SALT, we refine these weights using the following procedure. First, consider a weight matrix $\mathbf{W}^{(l)}$. SALT decomposes $\mathbf{W}^{(l)}$ using Singular Value Decomposition (SVD).
$\mathbf{W}^{(l)}=\mathbf{U}^{(l)}\,\mathbf{\Sigma}^{(l)}\,\bigl{(}\mathbf{V}^{(l)}\bigr{)}^{\top},$ (2)
where $\mathbf{U}^{(l)}\in\mathbb{R}^{D\times D}$ (left singular vectors), $(\mathbf{V}^{(l)})^{\top}\in\mathbb{R}^{K\times K}$ (right singular vectors), and $\mathbf{\Sigma}^{(l)}\in\mathbb{R}^{D\times K}$ is a diagonal matrix of the singular values corresponding to each singular vector. We denote $min(D,K)$ as $r_{max}$ and split this matrix into two matrices (i) $\mathbf{\Sigma}^{(l)}_{r}$ which contains a diagonal vector of top r singular values where $r\leq r_{max}$ (ii) $\mathbf{\Sigma}^{(l)}_{r^{\prime}}$ containing the remaining $r^{\prime}$ singular values where $r^{\prime}=r_{max}-r$. Then SALT applies (i) trainable scale and shift transformations to the top $r$ singular values. Concretely, define parameters $\alpha^{(l)}$ and $\beta^{(l)}$ as a diagonal matrices $\in\mathbb{R}^{r\times r}$ to transform $\mathbf{\Sigma}^{(l)}_{r}$:
$\mathbf{\Sigma}^{{}^{\prime}(l)}_{r}=\alpha^{(l)}\odot\mathbf{\Sigma}^{(l)}_{r}+\beta^{(l)},$ (3)
where $\odot$ denotes Hadamard product, and (ii) a LoRA-based low-rank update $\mathbf{X}^{(l)}\mathbf{Y}^{(l)\top}$ where $\mathbf{X}^{(l)}\in\mathbb{R}^{r^{\prime}\times d_{\text{lora}}}$ , $\mathbf{Y}^{(l)}\in\mathbb{R}^{d_{\text{lora}}\times r^{\prime}}$ to the remaining singular values $\mathbf{\Sigma}^{(l)}_{r^{\prime}}$. The updated modified singular matrix $\mathbf{\Sigma}^{{}^{\prime}(l)}$ is then activated elementwise by $\mathrm{ReLU}$ to preserve semi-definiteness, yielding the updated weight matrix:
@@
$\widetilde{\mathbf{W}}^{(l)}\;=\;\mathbf{U}^{(l)}\,\bigl{[}\mathrm{ReLU}\!\bigl{(}{\mathbf{\Sigma}}^{{}^{\prime}(l)}\bigr{)}\bigr{]}\,\bigl{(}\mathbf{V}^{(l)}\bigr{)}^{\top}.$ (4)
@@
During fine-tuning, only the newly introduced SALT parameters $\{\alpha^{(l)},\beta^{(l)},\mathbf{X}^{(l)}$, $\mathbf{Y}^{(l)}\}$ are optimized along with Normalization Layers and Text Affine Layers, while all original SAM weights remain frozen. The prompt encoder $\mathcal{E}_{\mathrm{p}}$, receiving text prompts $\mathbf{p}$, yields additional embeddings merged with the image-encoder $\mathcal{E}_{\mathrm{i}}$ output in the mask decoder $\mathcal{D}_{\mathrm{m}}$. We define SALT, denoted $S^{\prime}$, as the combination of two components: The original SAM model’s frozen output $S(i;\Theta_{\text{fixed}})$, which remains static during fine-tuning, and a lightweight, learnable adjustment term $\Delta S(i;\theta)$, introduced by SALT layers,capturing task-specific refinements through the trainable parameters $\theta$. This formulation ensures that the core SAM backbone remains unchanged, while task-specific adaptations are layered on top via the SALT modules. Putting this together, our objective is to minimize the loss function $\mathcal{L}(S^{\prime}(i),y)$ where $i$ is an input sample and $y$ is the corresponding ground truth label. To train SALT, we use a combination of segmentation losses, including a Focal loss, a Dice loss, and a regularization term expressed as:
$\mathcal{L}_{\text{reg}}=\underbrace{\|\alpha\odot s+\beta-s\|_{F}}_{\text{Scale-Shift}}+\underbrace{\|XY\|_{F}}_{\text{LoRA}},$ (5)
@@
where $\|\cdot\|_{F}$ denotes Frobenius norm. This dual constraint balances two goals: retaining the useful, pre-trained features and making only the minimal modifications needed for effective domain adaptation.
@@
Rationale for the SALT Design As shown in Fig. 2, LoRA [7] modifies a weight matrix $W\in\mathbb{R}^{D\times K}$ by adding $XY$, where $X\in\mathbb{R}^{D\times r}$ and $Y\in\mathbb{R}^{r\times K}$ have rank $r\ll\min(D,K)$. This reduces trainable parameters from $D\times K$, down to $r(D+K)$, but can underfit if $r$ is too small. By contrast, SVD-based fine tuning in S-SAM [16], shown in Fig. 2, learns only the scale and shift of each singular value in $W$. Since $\mathrm{rank}(W)=\min(D,K)$, this corresponds to $2\times\min(D,K)$ parameters, fewer than LoRA, yet fully updating every singular value may lack flexibility for subtle domain shifts. SALT in Fig. 2 merges the strengths of both. It trains the top singular values via scale and shift and uses LoRA blocks for the lower singular subspace. This dual strategy retains key features while capturing domain-specific nuances, balancing parameter efficiency with adaptation capability.

### v2

# Abu Dhabi, UAE SALT

@@ Abstract
The complex nature of medical image segmentation calls for models that are specifically designed to capture detailed, domain-specific features. Large foundation models offer considerable flexibility, yet the cost of fine-tuning these models remains a significant barrier. Parameter-Efficient Fine-Tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA), efficiently update model weights with low-rank matrices but may suffer from underfitting when the chosen rank is insufficient to capture domain-specific nuances. Conversely, full-rank Singular Value Decomposition (SVD) based methods provide comprehensive updates by modifying all singular values, yet they often lack flexibility and exhibit variable performance across datasets. We propose SALT (Singular Value Adaptation with Low-Rank Transformation), a method that selectively adapts the most influential singular values using trainable scale and shift parameters while complementing this with a low-rank update for the remaining subspace. This hybrid approach harnesses the advantages of both LoRA and SVD, enabling effective adaptation without relying on increasing the size or depth of the model. Evaluated on 5 challenging medical datasets, ranging from as few as 20 samples to 1000, SALT outperforms state-of-the-art PEFT (LoRA and SVD) by 2% to 5% in Dice with only 3.9% trainable parameters, demonstrating robust adaptation even in low-resource settings. The code for this paper is available at : https://github.com/BioMedIA-MBZUAI/SALT.
footnotetext: * Equal contribution
@@ 1 Introduction
Medical image segmentation plays a pivotal role in modern healthcare by enabling precise delineation of anatomical structures, pathological regions, and functional tissues across diverse imaging modalities such as CT, MRI, X-ray, ultrasound, and endoscopy [Litjens et al.(2017)Litjens, Kooi, Bejnordi, Setio, Ciompi, Ghafoorian, van der Laak, van Ginneken, and Sánchez, Ronneberger et al.(2015)Ronneberger, Fischer, and Brox]. While Deep Learning (DL) has revolutionized this field, with architectures like U-Net [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox] and nnU-Net [Isensee et al.(2021)Isensee, Jaeger, Kohl, Petersen, and Maier-Hein] achieving state-of-the-art (SOTA) performance through sophisticated designs and extensive data augmentation, these models often rely on large parameter counts, limiting their adaptability to new datasets or modalities. For instance, nnU-Net’s success hinges on dynamic architecture adjustments and aggressive augmentation pipelines, while DeepLabv3+ [Chen et al.(2018)Chen, Zhu, Papandreou, Schroff, and Adam] employs computationally intensive components like spatial pyramid pooling. Recent advances in foundation models have introduced specialized tools for diverse computer vision tasks. CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever] and ALIGN [Jia et al.(2021)Jia, Yang, Xia, Chen, Parekh, Pham, Le, Sung, Li, and Duerig], trained on large-scale image-text datasets, excel at cross-modal understanding for zero-shot image classification and open-set retrieval. In contrast, the Segment Anything Model (SAM) [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollar, and Girshick, Ravi et al.(2024)Ravi, Gabeur, Hu, et al.], trained on billions of natural images, exhibits robust zero-shot generalization capabilities and initially raised hopes for universal segmentation. However, SAM’s performance on medical data has proven suboptimal, with studies [Zhang et al.(2024)Zhang, Shen, and Jiao, He et al.(2023)He, Bao, Li, Stout, Bjornerud, Grant, and Ou] demonstrating its inferiority to specialized models like U-Net++ [Zhou et al.(2018)Zhou, Rahman Siddiquee, Tajbakhsh, and Liang], particularly in Dice scores across 12 medical datasets. Even with prompt engineering, SAM’s zero-shot accuracy remains below fully supervised methods [Cheng et al.(2023)Cheng, Qin, Jiang, Zhang, Lao, and Li], underscoring the domain gap between natural and medical images.
@@
Recent efforts to adapt SAM to medical imaging, such as MedSAM [Ma et al.(2024)Ma, He, Li, Han, You, and Wang], SAM Adapter [Wu et al.(2023)Wu, Fu, Fang, Liu, Wang, Xu, Jin, and Arbel], and S-SAM [Paranjape et al.(2024)Paranjape, Sikder, Vedula, and Patel], have made strides by incorporating medical training data, low-rank adapters, or SVD-based tuning. Yet challenges persist: PEFT methods must balance minimal parameter overhead, computational feasibility, and preservation of pre-trained knowledge while addressing unique medical imaging challenges like speckle noise, low contrast, and modality-specific artifacts. Promising approaches that arose to adapt foundational models, especially language models, like LoRA [Hu et al.(2022)Hu, yelong shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen] and AdaLoRA [Zhang et al.(2023)Zhang, Chen, Bukharin, He, Cheng, Chen, and Zhao] approximate weight updates via low-rank matrices but struggle to capture both dominant and nuanced features, while SVD-based methods like PiSSA [Meng et al.(2024)Meng, Wang, and Zhang] focus on principal singular vectors to capture the most significant data patterns, but this rigid prioritization can limit their adaptability to dynamic or fine-grained data characteristics. To address these limitations, we propose Singular Value Adaptation with Low-Rank Transformation (SALT), a novel PEFT framework that synergizes the strengths of SVD and low-rank adaptation. SALT selectively scales critical singular values to preserve foundational knowledge while applying trainable low-rank transformations to residual components, enabling efficient adaptation to medical domains with minimal parameter overhead. Evaluated on five challenging medical datasets, with sample sizes ranging from 20 to 1000 and demanding a high level of detailed segmentation, SALT consistently outperforms LoRA and SVD, achieving a 2% to 5% improvement in Dice score while utilizing only 3.9% trainable parameters. Our contributions are threefold:
  • We introduce SALT, a hybrid PEFT approach that adapts dominant singular values via SVD while applying low-rank residuals, enabling efficient domain adaptation.
@@
  • We demonstrate SALT’s utility through extensive experiments on diverse medical datasets from various modalities, showing improvements over SOTA PEFT methods.
@@ 2 Methodology
2.1 Problem Formulation
Let S be a SAM model with pretrained parameters $\Theta$, which consists of an image encoder ($\mathcal{E}_{\mathrm{i}}$), a prompt encoder ($\mathcal{E}_{\mathrm{p}}$), a mask decoder ($\mathcal{D}_{\mathrm{m}}$). As depicted in Fig. 1, $\mathcal{E}_{\mathrm{i}}$ contains $L$ blocks, each comprising a Multi-Head Attention module (MHA) and a Multi-Layer Perceptron (MLP). For a given input image $\mathbf{i}$, the encoder $\mathcal{E}_{\mathrm{i}}$ produces intermediate feature maps $\{\mathbf{f}^{(l)}\}_{l=1}^{L}$. Therefore, in the $l$-th block, the main operations can be written as:
$q,k,v=W_{qkv}^{l}\,x^{(l-1)}+b_{qkv}^{l},\quad o=W_{M}^{l}\,x^{(l-1)}+b_{M}^{l},$ (1)
where $x^{(l-1)}$ is the input to block $l$ (the output of block l1l\!-\!1), $q,k,v$ are the query, key and value projection tensors produced by the MHA module in block $l$ (that is, the inputs to the attention computation) and $W_{qkv}^{l},W_{M}^{l}\in\mathbb{R}^{D\times K}$ and $b_{qkv}^{l},b_{M}^{l}\in\mathbb{R}^{D}$ are the attention/MLP weights and biases for block $l\in\{1,...,L\}$. While effective for natural images, the pre-trained weights $W\in\mathbb{R}^{D\times K}$ may not capture all domain-specific nuances.
2.2 Singular Value Adaptation with Low-Rank Transformation
Through SALT, we refine these weights using the following procedure to obtain an adapted SAM model $S^{\prime}$, shown in Fig. 1. First, consider a weight matrix $\mathbf{W}^{(l)}$. SALT decomposes $\mathbf{W}^{(l)}$ using Singular Value Decomposition (SVD).
$\mathbf{W}^{(l)}=\mathbf{U}^{(l)}\,\boldsymbol{\Sigma}^{(l)}\,\bigl{(}\mathbf{V}^{(l)}\bigr{)}^{\top},$ (2)
where $\mathbf{U}^{(l)}\in\mathbb{R}^{D\times D}$ (left singular vectors), $(\mathbf{V}^{(l)})^{\top}\in\mathbb{R}^{K\times K}$ (right singular vectors), and $\boldsymbol{\Sigma}^{(l)}\in\mathbb{R}^{D\times K}$ collects the singular values on its main diagonal. We denote $min(D,K)$ as $r_{max}$ and split this matrix into two matrices (i) $\boldsymbol{\Sigma}^{(l)}_{r}$ which contains a diagonal vector of top r singular values where $r\leq r_{max}$ (ii) $\boldsymbol{\Sigma}^{(l)}_{r^{\prime}}$ containing the remaining $r^{\prime}$ singular values where $r^{\prime}=r_{max}-r$. Then SALT applies (i) trainable scale and shift transformations to the top $r$ singular values. Concretely, define parameters $\alpha^{(l)}$ and $\beta^{(l)}$ as diagonal matrices $\in\mathbb{R}^{r\times r}$ to transform $\boldsymbol{\Sigma}^{(l)}_{r}$:
$\boldsymbol{\Sigma}^{{}^{\prime}(l)}_{r}=\alpha^{(l)}\odot\boldsymbol{\Sigma}^{(l)}_{r}+\beta^{(l)},$ (3)
where $\odot$ denotes Hadamard product, and (ii) a LoRA-based low-rank update $\mathbf{X}^{(l)}\mathbf{Y}^{(l)\top}$ where $\mathbf{X}^{(l)}\in\mathbb{R}^{r^{\prime}\times d_{\text{lora}}}$ , $\mathbf{Y}^{(l)}\in\mathbb{R}^{d_{\text{lora}}\times r^{\prime}}$ to the remaining singular values $\boldsymbol{\Sigma}^{(l)}_{r^{\prime}}$. The updated core matrix $\boldsymbol{\Sigma}^{{}^{\prime}(l)}$ is then activated elementwise by $\mathrm{ReLU}$ to ensure nonnegative entries, yielding the updated weight matrix:
@@ 2.2 Singular Value Adaptation with Low-Rank Transformation
$\widetilde{\mathbf{W}}^{(l)}\;=\;\mathbf{U}^{(l)}\,\bigl{[}\mathrm{ReLU}\!\bigl{(}{\boldsymbol{\Sigma}}^{{}^{\prime}(l)}\bigr{)}\bigr{]}\,\bigl{(}\mathbf{V}^{(l)}\bigr{)}^{\top}.$ (4)
@@
During fine-tuning, only the newly introduced SALT parameters $\{\alpha^{(l)},\beta^{(l)},\mathbf{X}^{(l)}$, $\mathbf{Y}^{(l)}\}$ are optimized along with the normalization layers and the Text Affine Layer (TAL), while all original SAM weights remain frozen. Unlike the SVD’s $\Sigma^{(l)}$, $\Sigma^{\prime(l)}$ is not necessarily diagonal due to the added $\mathbf{X}^{(l)}\mathbf{Y}^{(l)\top}$ block. Embeddings of the text prompts are acquired using a frozen CLIP model [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever], after which TAL projects the CLIP embeddings to the mask decoder ($\mathcal{D}_{\mathrm{m}}$) token space [Paranjape et al.(2024)Paranjape, Sikder, Vedula, and Patel]. The prompt encoder $\mathcal{E}_{\mathrm{p}}$, processes the projected CLIP embeddings $p$, and merges them with the image-encoder $\mathcal{E}_{\mathrm{i}}$ output in the mask decoder $\mathcal{D}_{\mathrm{m}}$. We define SALT, denoted $S^{\prime}$, as the combination of two components: The original SAM model’s frozen output $S(i;\Theta_{\text{fixed}})$, which remains static during fine-tuning, and a lightweight, learnable adjustment term $\Delta S(i;\theta)$, introduced by SALT layers, capturing task-specific refinements through the trainable parameters $\theta$. This formulation ensures that the core SAM backbone remains unchanged, while task-specific adaptations are layered on top via the SALT modules. Putting this together, our objective is to minimize the loss function $\mathcal{L}(S^{\prime}(i),y)$ where $i$ is an input sample and $y$ is the corresponding ground truth label.
SALT performs the SVD factorization only once per layer at initialization; after that, all training steps update just the low-dimensional scale, shift, and LoRA parameters, keeping the computational overhead low. To train SALT, we use a combination of segmentation losses, including a Focal loss, a Dice loss, and a regularization term expressed as:
$\mathcal{L}_{\mathrm{reg}}=\sum_{l=1}^{L}\Big{\|}\alpha^{(l)}\odot\boldsymbol{\Sigma}^{(l)}_{r}+\beta^{(l)}-\boldsymbol{\Sigma}^{(l)}_{r}\Big{\|}_{F}\;+\;\big{\|}\mathbf{X}^{(l)}\mathbf{Y}^{(l)}\big{\|}_{F}$ (5)
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm. This dual constraint balances two goals: retaining the useful, pre-trained features and making only the minimal modifications needed for effective domain adaptation.
@@
Rationale for the SALT Design As shown in Fig. 2, LoRA [Hu et al.(2022)Hu, yelong shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen] modifies a weight matrix $W\in\mathbb{R}^{D\times K}$ by adding $XY$, where $X\in\mathbb{R}^{D\times r}$ and $Y\in\mathbb{R}^{r\times K}$ have rank $r\ll\min(D,K)$. This reduces trainable parameters from $D\times K$, down to $r(D+K)$, but can underfit if $r$ is too small. By contrast, SVD-based fine tuning in S-SAM [Paranjape et al.(2024)Paranjape, Sikder, Vedula, and Patel], shown in Fig. 2, learns only the scale and shift of each singular value in $W$. Since $\mathrm{rank}(W)=\min(D,K)$, this corresponds to $2\times\min(D,K)$ parameters, fewer than LoRA, yet fully updating every singular value may lack flexibility for subtle domain shifts. SALT in Fig. 2 merges the strengths of both. It trains the top singular values via scale and shift and uses LoRA blocks for the lower singular subspace. This dual strategy retains key features while capturing domain-specific nuances, balancing parameter efficiency with adaptation capability.